# Self-organized transport in noisy dynamic networks

### v1

@@ I Introduction
Figure 1: Network self-organization for a multi-commodity flow problem in the presence of stochastic fluctuations. The sources and sinks simulate the metro stations of Tokyo railroad on a grid of 31 x 31 nodes with 528 demands, as illustrated in (a). The red dots represent the relative locations of the major cities around Tokyo, Tokyo itself is indicated by the green dot. The network design results from the dynamics of the edges, which are modelled by time-varying edge capacities (Eq. (10)). The emerging network topology is shown in (b) for the noiseless case ($\alpha=0$). Subplot (c) displays a single trajectory in the presence of noise ($\alpha=0.002$) and for a specific choice of the activation function (sigmoidal with exponent $n=1.6$). The widths of the edges are proportional to the corresponding capacities. The backbone of the network in (c) is extracted using a filtering procedure, the filtered network is shown in (d). In all simulations, we initially set all edge capacities equal to the value $D_{u,v}^{0}=0.5$. Furthermore, for each commodity $i$, we fix the potential $p_{x}^{i}=0$ at the node indicated by the yellow cross. Further details on the numerical simulations are reported in the main text.
@@
In the present study we consider a multi-commodity flow problem in the presence of noise. We treat the edge capacity as a variable, that depends on the current flowing across the edge by means of a nonlinear (activation) function Tero et al. (2007), and further assume that the edge capacity can undergo random fluctuations. We analyse the response to noise for three different functional behaviors of the activation function. As in Ref. Tero et al. (2010); Bonifaci et al. (2022), we consider a graph with a number of demands geometrically arranged to reflect the relative locations of the major cities around Tokyo, which is illustrated in Fig. 1(a). We determine the characteristics of the emerging networks as a function of the noise amplitude and, in particular, analyse their robustness, transport efficiency, and cost. Figures 1(b) and (c) provide examples.
@@
This paper is organized as follows. In Sec. II, we introduce the model and define the measures used to evaluate the emerging network topologies. We then determine the network measures as a function of the demand strength for the three different activation functions with no noise. Section III is devoted to the numerical methods and the algorithms used to extract the network topology from the stochastic dynamics. The network measures as a function of the noise amplitude are presented in Sec. IV. Conclusions and outlook are drawn in Sec. V.
@@ II Noiseless transport on a grid
In this section, we introduce the model and the network measures used for evaluating the topologies. We focus on noiseless transport and study the network topology for different activation functions.
@@ II.1 The model
Two-norm $f(\{Q_{u,v}^{(j)}\})=\sqrt{(Q_{u,v}^{(1)})^{2}+...+(Q_{u,v}^{(k)})^{2}}$
Hill/sigmoidal $f(\{Q_{u,v}^{(j)}\})=(Q_{u,v})^{n}/(\kappa^{n}+(Q_{u,v})^{n})$ $\kappa=1$
ReLU $f(\{Q_{u,v}^{(j)}\})=(Q_{u,v}-\kappa)\theta(Q_{u,v}-\kappa)$ $\kappa=0.01$
Table 1: Three classes of activation functions. Here, $Q_{u,v}$ denotes the total flow along the edge, which is determined as $Q_{u,v}=\sum_{j}|Q_{u,v}^{(i)}|$ (one norm). For the Hill function, we consider $n=1.6$ and $n=2$. In the definition of the ReLU function, $\theta(x)$ is Heaviside’s function ($\theta(x)=1$ for $x>0$ and $\theta(x)=0$ otherwise).
We analyze network self-organization for the given multi-commodity flow problem represented by the distribution of sources and sinks, illustrated by the red circles in Fig. 1(a). Pairs of sources ($s_{+}^{i}$) and sinks ($s_{-}^{i}$) form a demand, where a flow is injected at the source ($+I_{i}$) and extracted at the sink ($-I_{i}$). We use the same setup as in Ref. Tero et al. (2010); Bonifaci et al. (2022), where the distribution of sources and sinks mimics the relative locations of the major cities around Tokyo. We assume that there is a transport demand between every pair of cities, meaning that the emerging network needs to serve $k=528$ demands.
The sources and sinks are geometrically arranged on a grid that discretizes the space and is composed of $N=31\times 31$ nodes, of which 33 are the cities. Each node, labelled $u=1,\ldots,N$, is connected to a number of nearest and next-nearest neighbors, described by the set $E_{u}$. We denote the edge connecting nodes $u$ and the neighbor $v\in E_{u}$ by the pair $(u,v)$. The edge length is $L_{u,v}$ and is set to unity when the nodes $u$ and $v$ are nearest neighbors and to $\sqrt{2}$ when the nodes are connected by a diagonal.
We choose $I_{i}=I_{0}$ for all demands $i$ except for those that involve the node at the relative location of Tokyo: For these demands, as in Ref. Tero et al. (2010), we set $I_{i}=7\,I_{0}$ to reflect the importance of Tokyo as the center of the region. The demand $i$ is realized by a flow $I_{i}$ across the edges of the network. The flow through the edge connecting nodes $u$ and $v$ is denoted as $Q_{u,v}^{i}$. At the sources and sinks, it obeys the constraints:
$\sum_{v\in E_{s_{i}^{\pm}}}Q_{s_{i}^{\pm},v}^{i}=\pm I_{i}\,.$ (1)
@@ II.2 The potential
The potential $p_{u}^{i}$ is determined for each demand $i$ as a function of $D_{u,v}(t)$ by solving the linear set of equations in Eq. (3) with Kirchhoff’s law. Let $N$ be the number of network nodes and $b^{i}\in\mathbb{R}$ be the vector representing the constraints, such that it has a value of $+I_{i}$ at the source node $s_{+}^{i}$ and $-I_{i}$ at the sink $s_{-}^{i}$. Otherwise, the entry is zero. The vector $p^{i}=(p_{1}^{i},\ldots,p_{N}^{i})$ contains all node potentials associated to demand $i$ and is determined by the linear system of equations Bonifaci et al. (2022); Lonardi et al. (2022):
@@ II.3 The edge capacity (conductivity)
$\partial_{t}D_{u,v}=f(\{Q_{u,v}^{(j)}\})-\gamma D_{u,v}\,,$ (6)
where $f(\{Q_{u,v}^{(j)}\})$ is the activation function modelling the feedback of the flows associated to the commodities $j=1,\ldots,k$ on the evolution of the edge conductivity. The edge capacities are damped with rate $\gamma$ and hence edges with little flow on them (in particular, flow below a fixed threshold) are suppressed.
The network topology is determined by the resulting edge capacities $D_{u,v}$. They are found by integrating Eq. (6) using Eq. (3), where the potential is found from Eq. (4) imposing the static boundary conditions, at the sources and sinks, Eq. (1), and Kirchhoff’s law otherwise.
@@ II.4 The activation function
Figure 2: Network measures for the noiseless dynamics and as a function of the demand strength $I_{0}$. The subplot display (a) the robustness $r$, (b) the transport efficiency $\sigma$ and (c) the cost $c$. They are extracted from the network obtained by integrating the coupled equations (6), (3) and (4) for a time of $1000\,\gamma^{-1}$ using a different activation function in each case, see legenda and Tab. 1. The measures of the networks are calculated after applying the disparity filter, which we detail in Sec. III.2. In the rest of this work, we will focus on the range of $I_{0}$ indicated by the shaded regions.
@@
The sigmoidal function saturates: This limits the flow that can be transported through the edge. As the flow between two nodes approaches the maximum capacity of the edge, the dynamics tends to construct multiple connections between them. In the following, we consider two different powers $n$ for the Hill function and analyse the regime where the dynamics is sensitive to small gradients as the demand strength $I_{0}$ increases above threshold.
@@ II.5 Performance measures of networks
$R_{i}=(p_{s_{+}^{i}}^{i}-p_{s_{-}^{i}}^{i})/I_{i}$
is the effective resistance between the source node $s_{+}^{i}$ and the sink node $s_{-}^{i}$ of demand $i$. The quantity $R_{i}$ takes into account both the number of connections between $s_{+}^{i}$ and $s_{-}^{i}$ as well as their lengths. The measure $r$ is an indicator of robustness against edge failure. In fact, adding edges and/or reducing the length of connections increases the measure Ellens et al. (2011). This is a different, nevertheless similar approach to existing measures, where the fault-tolerance of a network is measured by counting the number of edges that can be removed without separating the network into two parts for example Tero et al. (2010). Note that we set all non-vanishing edge conductivities to the same value (here, $D_{u,v}=0.5$) before calculating $r$, corresponding to a fixed resistance per length.
The transport efficiency $\sigma$ of the network is defined as the average over all demands of the length of the shortest path $d_{i}$ connecting the source and the sink node $s_{i}^{+}$ and $s_{i}^{-}$. It reads
$\sigma=\frac{k}{\sum_{i=1}^{k}d_{i}}\,.$ (8)
@@
Figure 3: Network measures as a function of the demand strength $I_{0}$ for the two-norm function and in the presence of Gaussian noise ($\alpha=0.001$). The panel displays (a) the robustness $r$, (b) the transport efficiency $\sigma$, and (c) the cost $c$. The color scale at a given value of $I_{0}$ indicates the percentage of 250 trajectories at the corresponding value of the measure (dark blue is statistically irrelevant, dark red corresponds to more than 20%). The solid white line indicates the average for each value of $I_{0}$. The solid purple line shows the deterministic value ($\alpha=0$) for comparison. The network measures are calculated after applying the disparity filter (see text). Details on the numerical simulations are reported in Sec. III.1.
@@ III Model and numerical methods for the stochastic dynamics
$\partial_{t}D_{u,v}=f(\{Q_{u,v}^{(j)}\})-\gamma D_{u,v}+\sqrt{\gamma}\alpha\xi_{u,v}(t)$ (10)
with the stochastic force $\xi_{u,v}(t)$, whose amplitude is scaled by the parameter $\alpha$. The force is statistically defined by the average over an ensemble of trajectories: it has no net drift, $\langle\xi_{u,v}(t)\rangle=0$, and simulates Gaussian white noise with no spatial correlations, $\langle\xi_{u,v}(t)\xi_{u^{\prime},v^{\prime}}(t^{\prime})\rangle=\delta_{u,u^{\prime}}\delta_{v,v^{\prime}}\delta(t-t^{\prime})$ VAN KAMPEN (2007). The network dynamics results from integrating the coupled equations (3), (4) and (10).
@@ III.1 Numerical methods
The integration of Eq. (10) is performed using the Euler-Maruyama scheme, as outlined in Ref. Kloeden and Platen (1992). We have analyzed convergence for different step sizes: For the parameter values considered here, we set the step size $\Delta t=0.05\,\gamma^{-1}$. We then set the evolution time as $t_{\rm end}=1000\,\gamma^{-1}$. In fact, extensive tests over a statistically relevant ensemble of trajectories (namely, individual network evolutions) show that each trajectory has reached a (meta-)stable configuration. In all simulations, the conductivities of all edges are initially equal to the value $D_{u,v}^{0}=0.5$. For each commodity $i$, we fix the potential $p_{x}^{i}=0$ at the node indicated by the cross in Fig. 1(a).
@@ IV Network topologies in the presence of noise
Figure 5: The solid line indicates the difference $\Delta r=\langle r\rangle-r_{0}$ between the ensemble average $\langle r\rangle$ of the robustness and the value $r_{0}$ of the noiseless dynamics as a function of the noise amplitude $\alpha$ for $I_{0}=4.5\times 10^{-2}$. Each data point represents the average value over 250 simulation runs. The red dots indicate the difference $\Delta r$ after applying the deterministic filter to the noisy networks for a time $t=2000\,\gamma^{-1}$.
@@
An important result is that the topology depends on $I_{0}$. This is in striking contrast with the noiseless behavior (magenta curves), where all measures are independent of the demand strength $I_{0}$. The average robustness, transport efficiency, and cost are generally increasing as a function for $I_{0}$. For small $I_{0}$ ($I_{0}\sim 1.5\times 10^{-2}$), the mean values of the robustness, transport efficiency and cost are lower than in the noiseless case. For larger values of $I_{0}$ ($I_{0}\sim 5.5\times 10^{-2}$), the mean value of robustness and transport efficiency slightly exceed the corresponding noiseless measures (see Figs. 3(a) and (b)), whereas the average cost is smaller. At values of $I_{0}$ outside the interval, these measures all saturate to their values in the noiseless case. Astonishingly, there is a regime where the dynamics converges to more efficient network designs, namely, networks that are more robust and efficient, and at the same time less costly.
@@
It is tempting to assume that noise will lead to inferior networks. This is true for large enough noise. However, small noise can be beneficial. In the dark red part of the diagram the noisy solutions are more robust and efficient than the noiseless networks and the network topologies reach the largest values of $r$ and $\sigma$. A cut of the plot at a fixed $I_{0}$, Fig. 5, indicates a resonance-like behavior as a function of the noise amplitude $\alpha$. Remarkably, in this region, the standard deviation of the robustness decreases, see Fig. 6. This behavior challenges the conventional expectation that the size of the fluctuations increase with $\alpha$.
Figure 6: The standard deviation of the robustness distribution shown in Fig. 4 as a function of the demand strength $I_{0}$ and the noise amplitude $\alpha$.
Figure 7: Network measures as a function of the demand strength $I_{0}$ and for $\alpha=0.001$. The panels in row (a) are obtained using the sigmoidal function with $n=1.6$, in row (b) using the sigmoidal function with $n=2$, and in row (c) using the ReLU function. The panels display (left) the robustness $r$, (central) the transport efficiency $\sigma$ and (right) the cost $c$. The color scale at a given value of $I_{0}$ indicates the percentage of 250 trajectories at the corresponding value of $r$ (dark blue is statistically irrelevant, dark red corresponds to more than 20%). The solid white line indicates the average for each value of $I_{0}$. The solid purple line shows the noiseless value and is plotted for comparison. The results of the noisy dynamics are invariant under the filtering procedure we apply. Details on the numerical simulations are reported in Sec. III.1.
Following the criterion introduced in Ref. Folz et al. (2023), we denote the region of increased robustness and efficiency as a noise-induced resonance. This terminology is borrowed from nonlinear dynamics, where noise-induced coherent effects have been reported in models simulating forest fires Meron (1992) and neurons Lindner et al. (2004). As in stochastic resonance Gammaitoni et al. (1998); Lindner et al. (2004); Perc (2005); Folz et al. (2021), we observe an optimal range of noise amplitudes at which noise increases the efficiency of the algorithm leading to more robust network topology than in the noiseless case. We note that these solutions are not fixed points of the noiseless dynamics, as we verify by taking them as initial conditions and determine the evolution according to Eq. (6). This procedure, which we refer to as the deterministic filter and detail in Appendix A, shows that for sufficiently long integration times, $\Delta r\to 0$, as demonstrated by the red dots in Fig. 5. This suggests that the topologies obtained using the noisy dynamics are novel solutions, which are absent in the noiseless case.

### v2

@@ I Introduction
Figure 1: Network self-organization for a multi-commodity flow problem in the presence of stochastic fluctuations. The sources and sinks simulate the metro stations of Tokyo railroad on a grid of 31 x 31 nodes with 528 demands, as illustrated in (a). The red dots represent the relative locations of the major cities around Tokyo, Tokyo itself is indicated by the green dot. The network design results from the dynamics of the edges, which are modelled by time-varying edge capacities (Eq. (10)). The emerging network topology is shown in (b) for the noiseless case ($\alpha=0$). Subplot (c) displays a single trajectory in the presence of noise ($\alpha=0.002$) and for a specific choice of the activation function (sigmoidal with exponent $n=1.6$). The widths of the edges are proportional to the corresponding capacities. The backbone of the network in (c) is extracted using a filtering procedure, the filtered network is shown in (d). In all simulations, we initially set all edge capacities equal to the value $D_{u,v}^{0}=0.5$. Furthermore, for each commodity $i$, we fix the potential $p_{x}^{i}=0$ at the node indicated by the yellow cross as a reference. Further details on the numerical simulations are reported in the main text.
@@
In the present study, we analyze the influence of noise on multi-commodity flow problems, where the edge capacity is a variable that depends on the current flowing across the edge through a nonlinear (activation) function Tero et al. (2007). We assume that the edge capacity can undergo stochastic fluctuations and examine the emerging network topologies for three different functional behaviors of the activation function. In particular, we determine the characteristics of the emerging networks as a function of the noise amplitude and analyze their robustness, transport efficiency, and cost. For our analysis, we choose the same case study as in Refs. Tero et al. (2010); Bonifaci et al. (2022), where the graph has a number of demands geometrically arranged to mimic the relative locations of the major cities around Tokyo. The graph and the demands are illustrated in Fig. 1(a). Examples of the emerging networks are provided in Figs. 1(b) and (c).
@@
This paper is organized as follows. In Sec. II, we introduce the model and define the measures used to evaluate the emerging network topologies for a graph with demands mimicking the Tokyo railroad transport problem. We then determine the network measures as a function of the demand strength for the three different activation functions with no noise. Section III is devoted to the numerical methods and the algorithms used to extract the network topology from the stochastic dynamics. The network measures as a function of the noise amplitude are presented in Sec. IV. Conclusions and outlook are drawn in Sec. V.
@@ II Noiseless transport on a grid
In this section, we introduce the model, define the network measures used for evaluating the topologies and apply the multi-commodity flow problem to routing transport demands with the geometry of the greater area of Tokyo. We focus on noiseless transport and study the network topologies for different activation functions.
@@ II.1 The model
Two-norm $f(\{Q_{u,v}^{(i)}\})=\sqrt{(Q_{u,v}^{(1)})^{2}+...+(Q_{u,v}^{(k)})^{2}}$
Hill/sigmoidal $f(\{Q_{u,v}^{(i)}\})=(Q_{u,v})^{n}/(\kappa^{n}+(Q_{u,v})^{n})$ $\kappa=1$
ReLU $f(\{Q_{u,v}^{(i)}\})=(Q_{u,v}-\kappa)\theta(Q_{u,v}-\kappa)$ $\kappa=0.01$
Table 1: Three classes of activation functions. Here, $Q_{u,v}$ denotes the total flow along the edge, which is determined as $Q_{u,v}=\sum_{i}|Q_{u,v}^{(i)}|$ (one norm). For the Hill function, we consider $n=1.6$ and $n=2$. In the definition of the ReLU function, $\theta(x)$ is Heaviside’s function ($\theta(x)=1$ for $x>0$ and $\theta(x)=0$ otherwise).
The multi-commodity flow problem is represented by a set of $k$ transport demands on a graph that discretizes the space and is composed of $N$ nodes. Each node, labelled $u=1,\ldots,N$, is connected to a number of other nodes, described by the set $E_{u}$. We denote the edge connecting nodes $u$ and the neighbor $v\in E_{u}$ by the pair $(u,v)$. The edge length is given by $L_{u,v}$.
Each demand is formed by a pair of a source node $s_{+}^{i}$, where a flow is injected ($+I^{i}$), and a sink node ($s_{-}^{i}$), where it is extracted ($-I^{i}$). The demand $i$ is realized by a flow $I^{i}$ across the edges of the network. The flow through the edge connecting nodes $u$ and $v$ is denoted as $Q_{u,v}^{i}$. At the sources and sinks, it obeys the constraints:
$\sum_{v\in E_{s_{\pm}^{i}}}Q_{s_{\pm}^{i},v}^{i}=\pm I^{i}\,.$ (1)
@@ II.2 The potential
The potential $p_{u}^{i}$ is determined for each demand $i$ as a function of $D_{u,v}(t)$ by solving the linear set of equations in Eq. (3), subject to the boundary conditions at the source and sink nodes given by Eq. (1), and Kirchhoff’s law at any other node, Eq. (2). We fix the potential $p_{x}^{i}=0$ at the node indicated by the yellow cross in Fig. 1(a) as a reference. Let $N$ be the number of network nodes and $b^{i}\in\mathbb{R}^{N}$ the vector representing the constraints, such that it has a value of $+I^{i}$ at the source node $s_{+}^{i}$ and $-I^{i}$ at the sink $s_{-}^{i}$. Otherwise, the entry is zero. Then, the vector $p^{i}=(p_{1}^{i},\ldots,p_{N}^{i})$ containing all node potentials associated to demand $i$ is determined by the linear system of equations Bonifaci et al. (2022); Lonardi et al. (2022):
@@ II.3 The edge capacity (conductivity)
$\partial_{t}D_{u,v}=f(\{Q_{u,v}^{(i)}\})-\gamma D_{u,v}\,,$ (6)
where $f(\{Q_{u,v}^{(i)}\})$ is the activation function modelling the feedback of the flows associated to the commodities $j=1,\ldots,k$ on the evolution of the edge conductivity. The edge capacities are damped with rate $\gamma$ and hence edges with little flow on them (in particular, flow below a fixed threshold) are suppressed.
The network topology is determined by the resulting edge capacities $D_{u,v}$. They are found by integrating Eq. (6) using Eq. (3), where the potential is found from Eqs. (4) and (5).
@@ II.4 The activation function
Figure 2: Network measures for the noiseless dynamics and as a function of the demand strength $I_{0}$. The subplot display (a) the robustness $r$, (b) the transport efficiency $\sigma$ and (c) the cost $c$. They are extracted from the network obtained by integrating the coupled equations (3) - (6) for a time of $1000\,\gamma^{-1}$ using a different activation function in each case, see legenda and Tab. 1. The measures of the networks are calculated after applying the disparity filter, which we detail in Sec. III.2. In the rest of this work, we will focus on the range of $I_{0}$ indicated by the shaded regions.
@@
The sigmoidal function saturates: This limits the flow that can be transported through the edge. As the flow between two nodes approaches the maximum capacity of the edge, the dynamics tends to construct multiple connections between them. In the following, we consider two different powers $n$ for the Hill function and analyze the regime where the dynamics is sensitive to small gradients as the demand strength $I_{0}$ increases above threshold.
@@ II.5 Performance measures of networks
$R^{i}=(p_{s_{+}^{i}}^{i}-p_{s_{-}^{i}}^{i})/I^{i}$
is the effective resistance between the source node $s_{+}^{i}$ and the sink node $s_{-}^{i}$ of demand $i$. The quantity $R^{i}$ takes into account both the number of connections between $s_{+}^{i}$ and $s_{-}^{i}$ as well as their lengths. The measure $r$ is an indicator of robustness against edge failure. In fact, adding edges and/or reducing the length of connections increases the measure Ellens et al. (2011). This is a different, nevertheless similar approach to existing measures, where the fault-tolerance of a network is measured by counting the number of edges that can be removed without separating the network into two parts for example Tero et al. (2010). Note that we set all non-vanishing edge conductivities to the same value (here, $D_{u,v}=0.5$) before calculating $r$, corresponding to a fixed resistance per length.
The transport efficiency $\sigma$ of the network is defined as the average over all demands of the length of the shortest path $d^{i}$ connecting the source and the sink node $s_{+}^{i}$ and $s_{-}^{i}$. It reads
$\sigma=\frac{k}{\sum_{i=1}^{k}d^{i}}\,.$ (8)
@@
Figure 3: Network measures as a function of the demand strength $I_{0}$ for the two-norm function and in the presence of Gaussian noise ($\alpha=0.001$). The panel displays (a) the robustness $r$, (b) the transport efficiency $\sigma$, and (c) the cost $c$. The color scale at a given value of $I_{0}$ indicates the percentage $\Pi$ of 250 trajectories at the corresponding value of the measure (dark blue is statistically irrelevant, dark red corresponds to more than 20%). The solid white line indicates the average for each value of $I_{0}$. The solid purple line shows the deterministic value ($\alpha=0$) for comparison. The network measures are calculated after applying the disparity filter (see text). Details on the numerical simulations are reported in Sec. III.1.
@@
II.6 Application to the Tokyo transport problem
We apply the model introduced in this section to the multi-commodity flow problem represented by the distribution of sources and sinks illustrated by the red circles in Fig. 1(a). We use the same setup as in Refs. Tero et al. (2010); Bonifaci et al. (2022), where the sources and sinks are geometrically arranged to mimic the relative locations of the major cities around Tokyo. A transport demand between each pair of cities is assumed, which means that the emerging network has to satisfy $k=528$ demands.
We consider a grid of $N=31\times 31$ nodes, of which 33 are the cities. Each node, labelled $u=1,\ldots,N$, is connected to a number of nearest and next-nearest neighbors. The edge length $L_{u,v}$ is set to unity when the nodes $u$ and $v$ are nearest neighbors and to $\sqrt{2}$ when the nodes are connected by a diagonal.
We choose $I^{i}=I_{0}$ for all demands $i$ except for those that involve the node at the relative location of Tokyo: For these demands, as in Ref. Tero et al. (2010), we set $I^{i}=7\,I_{0}$ to reflect the importance of Tokyo as the center of the region.
@@ III Model and numerical methods for the stochastic dynamics
$\partial_{t}D_{u,v}=f(\{Q_{u,v}^{(i)}\})-\gamma D_{u,v}+\sqrt{\gamma}\alpha\xi_{u,v}(t)$ (10)
with the stochastic force $\xi_{u,v}(t)$, whose amplitude is scaled by the parameter $\alpha$. The force is statistically defined by the average over an ensemble of trajectories: it has no net drift, $\langle\xi_{u,v}(t)\rangle=0$, and simulates Gaussian white noise with no spatial correlations, $\langle\xi_{u,v}(t)\xi_{u^{\prime},v^{\prime}}(t^{\prime})\rangle=\delta_{u,u^{\prime}}\delta_{v,v^{\prime}}\delta(t-t^{\prime})$ VAN KAMPEN (2007). The network dynamics results from integrating the coupled equations (3), (4), (5) and (10).
@@ III.1 Numerical methods
The integration of Eq. (10) is performed using the Euler-Maruyama scheme, as outlined in Ref. Kloeden and Platen (1992). We have analyzed convergence for different step sizes: For the parameter values considered here, we set the step size $\Delta t=0.05\,\gamma^{-1}$. We then set the evolution time as $t_{\rm end}=1000\,\gamma^{-1}$. In fact, extensive tests over a statistically relevant ensemble of trajectories (namely, individual network evolutions) show that each trajectory has reached a (meta-)stable configuration. In all simulations, the conductivities of all edges are initially equal to the value $D_{u,v}^{0}=0.5$. For each commodity $i$, we fix the potential $p_{x}^{i}=0$ at the node indicated by the yellow cross in Fig. 1(a) as a reference.
@@ IV Network topologies in the presence of noise
Figure 5: (a) The solid line indicates the difference $\Delta r=\langle r\rangle-r_{0}$ between the ensemble average $\langle r\rangle$ of the robustness and the value $r_{0}$ of the noiseless dynamics as a function of the noise amplitude $\alpha$ for $I_{0}=4.5\times 10^{-2}$. Each data point represents the average value over 250 simulation runs. The red dots indicate the difference $\Delta r$ after applying the deterministic filter to the noisy networks for a time $t=2000\,\gamma^{-1}$. (b) The standard deviation of the robustness distribution shown in Fig. 4 as a function of the demand strength $I_{0}$ and the noise amplitude $\alpha$.
@@
We note that the topology depends on $I_{0}$. This is in striking contrast with the noiseless behavior (magenta curves), where all measures are independent of the demand strength $I_{0}$. The average robustness, transport efficiency, and cost are generally increasing as a function for $I_{0}$. For small $I_{0}$ ($I_{0}\sim 1.5\times 10^{-2}$), the mean values of the robustness, transport efficiency and cost are lower than in the noiseless case. For larger values of $I_{0}$ ($I_{0}\sim 5.5\times 10^{-2}$), the mean value of robustness and transport efficiency slightly exceed the corresponding noiseless measures (see Figs. 3(a) and (b)), whereas the average cost is smaller. At values of $I_{0}$ outside the interval, these measures all saturate to their values in the noiseless case. Remarkably, there is a regime where the dynamics converges to more efficient network designs, namely, networks that are more robust and efficient, and at the same time less costly.
@@
Figure 6: Network measures as a function of the demand strength $I_{0}$ and for $\alpha=0.001$. The panels in row (a) are obtained using the sigmoidal function with $n=1.6$, in row (b) using the sigmoidal function with $n=2$, and in row (c) using the ReLU function. The panels display (left) the robustness $r$, (central) the transport efficiency $\sigma$ and (right) the cost $c$. The color scale at a given value of $I_{0}$ indicates the percentage $\Pi$ of 250 trajectories at the corresponding value of $r$ (dark blue is statistically irrelevant, dark red corresponds to more than 20%). The solid white line indicates the average for each value of $I_{0}$. The solid purple line shows the noiseless value and is plotted for comparison. The results of the noisy dynamics are invariant under the filtering procedure we apply. Details on the numerical simulations are reported in Sec. III.1.
It is tempting to assume that noise will lead to inferior networks. This is true for large enough noise. However, small noise can be beneficial. In the dark red part of the diagram the noisy solutions are more robust and efficient than the noiseless networks and the network topologies reach the largest values of $r$ and $\sigma$. A cut of the plot at a fixed $I_{0}$, Fig. 5(a), indicates a resonance-like behavior as a function of the noise amplitude $\alpha$. Remarkably, in this region, the standard deviation of the robustness decreases, see Fig. 5(b). This behavior challenges the conventional expectation that the size of the fluctuations increase with $\alpha$.
Following the criterion introduced in Ref. Folz et al. (2023), we denote the region of increased robustness and efficiency as a noise-induced resonance. This terminology is borrowed from nonlinear dynamics, where noise-induced coherent effects have been reported in models simulating forest fires Meron (1992) and neurons Lindner et al. (2004). As in stochastic resonance Gammaitoni et al. (1998); Lindner et al. (2004); Perc (2005); Folz et al. (2021), we observe an optimal range of noise amplitudes at which noise increases the efficiency of the algorithm leading to more robust network topology than in the noiseless case. We note that these solutions are not fixed points of the noiseless dynamics, as we verify by taking them as initial conditions and determine the evolution according to Eq. (6). This procedure, which we refer to as the deterministic filter and detail in Appendix A, shows that for sufficiently long integration times, $\Delta r\to 0$, as demonstrated by the red dots in Fig. 5(a). This suggests that the topologies obtained using the noisy dynamics are novel solutions, which are absent in the noiseless case.